4. Suppose we run a multiple regression, and for observation 5 we get that Y5 = ‒4.73 and

Y5 = ‒4.28. What is the residual for observation 5?
Answer #1

The difference between the observed value of the dependent variable (y) and the predicted value (ŷ) is called the residual (e).

Formula of residual is given as following types

Residual = Observed value - Predicted value.

e= y - ŷ

= -4.28 - ( -4.73)

= -4.28+4.73

= 0.45

Please like ??